Grade Upto college level General Physics

What is self induction and self inductance? Derive an expression for the self inductance of a long solenoid carrying current

palash singh
12 Years agoGrade Upto college level

1 Answer

Apoorva Arora
12 Years ago
Self induction is that phenomenon in which a change in electric current in a coil produces an induced emf in the coil itself.
Mathematical representation:
Self induced emf in a coil is directly proportional to the rate of change of electric current in the coil i.e.
E\propto -\frac{\Delta I}{\Delta T}
E= -L\frac{\Delta I}{\Delta T}
where L= self inductance of the coil
Self Inductance:
Self inductance of a coil is defined as the ratio of self induced emf to the rate of change of current in the coil
L= \frac{E}{\Delta I/\Delta T}
It depends upon the physical characteristics of the coil and its unit is henry.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty